For each pair of functions, find (f+g)(x), (f-g)(x), and (fg)(x)...Then evauluate each function at 2.
1.) f(x)= 5x-10, g(x)= 3x+7
3.) f(x)= -4x+1, g(x)= 6x+2

Answer by tutorcecilia(2144)

The set-up is:
For each pair of functions, find (f+g)(x), (f-g)(x), and (fg)(x)...Then evauluate each function at 2.
1.) f(x)= 5x-10, g(x)= 3x+7
(f+g)(x)
(5x-10)+(3x+7)= [combine line-terms and subtitute x=2]
.
(f-g)(x)
(5x-10)-(3x+7)= [combine line-terms and subtitute x=2]
.
(fg)(x)
(5x-10)*(3x+7)= [multiply and subtitute x=2]]
3.) f(x)= -4x+1, g(x)= 6x+2
(-4x+1) + (6x+2)[combine line-terms and subtitute x=2]
.
(-4x+1) 1 (6x+2)[combine line-terms and subtitute x=2]
.
(-4x+1) * (6x+2)[multiply and subtitute x=2]